### v2

@@ 4 Inf-convolution of $\Lambda\mathrm{VaR}$ and one monotone risk measure with heterogenous beliefs
    For an increasing function $u$, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)=-\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$, then
    $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\left(x+(1-\beta_{x})u(-\infty)+\int_{0}^{\beta_{x}}u\left(\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)-x\right)\mathrm{d}t\right),$
    where $u(-\infty)=\lim_{y\to-\infty}u(y)$ and $\beta_{x}=\frac{(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})}{\mathbb{P}(B_{2})}$;
  3. (iii)
    For $\Lambda_{1}\in\mathcal{H}$ with $10<\lambda_{1}^{-}\leqslant\lambda_{1}^{+}<$ 1, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)=-\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$, then
@@
We mention that the results in Corollary 4 extend the results in the literature in the sense that $\Lambda$ is not a constant and the second risk measure is not law-invariant under $\mathbb{Q}_{1}$. For instance, (i) of Corollary 4 exactly extends Theorem 5.3 of Wang and Wei (2018) ($\rho$ is a distortion risk measure which is law-invariant under $\mathbb{Q}_{1}$) in these two directions. It is worth mentioning that (iii) of Corollary 4 is a new result even for the case $\mathbb{Q}_{1}=\mathbb{Q}_{2}$.
Let us now consider the case $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$. Instead of finding a simplified version of Theorem 5 for this case, we apply Theorem 5 directly to some specific $\rho$. For $\alpha\in(0,1)$, let $\mathrm{LES}_{\alpha}^{\mathbb{Q}}(X)=\frac{1}{\alpha}\int_{1-\alpha}^{1}\mathrm{VaR}_{t}^{\mathbb{Q}}(X)\mathrm{d}t$, which represents the lower conditional tail expectation.
@@
Suppose $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ with $\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}_{1}}$ denoted by $\eta$. For $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1, we have the following conclusions.
@@
    For an increasing function $u$, we have
    $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\left(x+u(-\infty)\Lambda(x)\mathrm{ES}_{\Lambda(x)}^{\mathbb{Q}_{1}}(\eta)+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{\{U_{\eta}^{\mathbb{Q}_{1}}<1-\Lambda(x)\}})\right),$
@@
Note that in (i) of Corollary 5, if $\Lambda$ is a constant, then the result coincides with Theorem 5 of Embrechts et al. (2020) for $n=2$, where the semi-explicit formula is offered. For the case $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$, it might be difficult to find more explicit expression for the inf-convolution of $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}$ and $\rho_{g}^{\mathbb{Q}_{2}}$/$\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}$.
@@ 5 Inf-convolution of $\Lambda\mathrm{VaR}^{+}$ and a monotone risk measure
The inf-convolution of multiple $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs is a very challenging problem even if we suppose the beliefs are homogeneous; see the discussion for homogeneous beliefs in Liu (2024). One important feature of $\Lambda\mathrm{VaR}^{+}$ is to capture the tail risk for decreasing $\Lambda$; see Frittelli et al. (2014) and Hitaj et al. (2018). In this section, we focus on the inf-convolution of $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}$ and a monotone risk measure $\rho^{\mathbb{Q}_{2}}$, which includes the inf-convolution of two $\Lambda\mathrm{VaR}^{+}$ as a special case. Let $\mathcal{X}=L^{\infty}$. For any $\Lambda\in\mathcal{H}$, we denote $\Lambda_{x}(z)=\inf_{x\leqslant t\leqslant z}\Lambda(t),~{}z\geqslant x$. Then $\Lambda_{x}(z)$ is decreasing and right-continuous for $z\geqslant x$. If $\Lambda$ is decreasing and right-continuous, then $\Lambda_{x}(z)=\Lambda(z),~{}~{}z\geqslant x$; and if $\Lambda$ is increasing, then $\Lambda_{x}(z)=\Lambda(x),~{}~{}z\geqslant x$. The following result is valid for general $\Lambda\in\mathcal{H}$.
@@
Suppose $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1, and $\rho$ is monotone and law-invariant under $\mathbb{Q}_{2}$. Then we have
@@
Moreover, the optimal allocation of the inf-convolution exists if and only if the minimizer of (9) exists. If $(x^{*},y^{*},U^{*})$ is the minimizer, then one optimal allocation is given by
@@
In the case of $\rho=\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}$ for $\Lambda_{1}\in\mathcal{H}$ with $10<\lambda_{1}^{-}\leqslant\lambda_{1}^{+}<$ 1, it follows from Theorem 6 that the optimal risk allocation for risk sharing between $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}$ and $\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}$ has the form of (10), which is neither comonotonic nor pairwise countermonotonic. This means that the optimal risk sharing for multiple $\Lambda\mathrm{VaR}^{+,\mathbb{Q}}$ is a very different problem from the cases with convex risk measures or quantile-based risk measures in the literature. The expression of the inf-convolution in (9) is an optimization problem involving three parameters $(x,y,U)$, where $U\sim U[0,1]$ under $\mathbb{Q}_{1}$. Note that the distribution of $U$ under $\mathbb{Q}_{2}$ depends on the relation between $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$. Moreover, the expression in (9) and also the optimal allocation in (10) show that shape of $\Lambda$ function plays an important role in both the expression of the inf-convolution and the form of the optimal allocations, which is very different from the results of $\Lambda\mathrm{VaR}$ displayed in Sections 3 and 4.
@@
For $x_{1}\in\mathbb{R}$ and $10<\lambda_{1}<\lambda_{2}<$ 1, let $\Lambda(z)=(1-\lambda_{1})\mathds{1}_{\{z<x_{1}\}}+(1-\lambda_{2})\mathds{1}_{\{z\geqslant x_{1}\}}$. Under the assumption of Theorem 6, we have
@@
where $\mathcal{B}_{\lambda_{1},\lambda_{2}}=\{(B_{1},B_{2}):\mathbb{Q}_{1}(B_{1})=\lambda_{1},\mathbb{Q}_{2}(B_{2})=\lambda_{2}-\lambda_{1},B_{1}\cap B_{2}=\emptyset\}$.
Note that the optimization problem in (11) is much easier than that in (9) because it only has parameters $(x,y,B_{1},B_{2})$.
@@ 6 Conclusion
In this paper, we consider the inf-convolution of multiple $\Lambda\mathrm{VaR}$ under heterogenous beliefs. We obtain the general expression for the inf-convolution and the corresponding optimal allocations. To study impact of the relation of the beliefs on the inf-convolution, we discuss three different cases: i) homogeneous beliefs; ii) conditional beliefs; iii) absolute continuous beliefs. For all these cases, we obtain more explicit expressions of the inf-convolution. In all above cases we demonstrate that trivial outcomes arise when both belief inconsistency and risk tolerance are high. The inf-convolution of one ΛVaR/$\Lambda\mathrm{VaR}/\Lambda\mathrm{VaR}^{+}$ and a general risk measures with belief heterogeneity are also discussed. There are still some unsolved problems such as the inf-convolution of multiple $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs, which deserves further study.
Acknowledgements: YW acknowledges financial support from the Natural Sciences and Engineering Research Council of Canada (RGPIN-2023-04674).
7 Declarations
Conflict of interest: All authors declare that they have no conflict of interest.
Appendix A Proof of Section 2

### v3

@@ 4 Inf-convolution of $\Lambda\mathrm{VaR}$ and one monotone risk measure with heterogeneous beliefs
    For $\Lambda_{1}\in\mathcal{H}$ with $0<\lambda_{1}^{-}\leqslant\lambda_{1}^{+}<1$, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)=$-$\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$, then
@@
We mention that the results in Corollary 5 extend the results in the literature in the sense that $\Lambda$ is not a constant and the second risk measure is not law-invariant under $\mathbb{Q}_{1}$. For instance, (i) of Corollary 5 exactly extends Theorem 5.3 of Wang and Wei (2018) ($\rho$ is a distortion risk measure which is law-invariant under $\mathbb{Q}_{1}$) in these two directions. It is worth mentioning that (ii) of Corollary 5 is a new result even for the case $\mathbb{Q}_{1}=\mathbb{Q}_{2}$.
Next, suppose $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$ has the relation (4). For an increasing function $u$, the expected utility is defined as $\mathbb{E}_{u}^{\mathbb{Q}}:X\mapsto\mathbb{E}^{\mathbb{Q}}(u(X))$. Instead of finding a simplified version of Theorem 5 for this case, we apply Theorem 5 directly to some specific $\rho$. For $\alpha\in(0,1)$, let $\mathrm{LES}_{\alpha}^{\mathbb{Q}}(X)=\frac{1}{\alpha}\int_{1-\alpha}^{1}\mathrm{VaR}_{t}^{\mathbb{Q}}(X)\mathrm{d}t$, which represents the lower conditional tail expectation.
@@
For $\Lambda\in\mathcal{H}$ with $0<\lambda^{-}\leqslant\lambda^{+}<1$, we have the following conclusions.
@@
    For $\mathcal{X}=L^{\infty}$ and an increasing function $u$, we have
    $=\inf_{x\in\mathbb{R}}\left(x+u(-\infty)(\mathbb{Q}_{2}(N)+\Lambda(x)\mathrm{ES}_{\Lambda(x)}^{\mathbb{Q}_{1}}(\eta))+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{\{U_{\eta}^{\mathbb{Q}_{1}}<1-\Lambda(x)\}})\right),$
@@
Note that in (i) of Corollary 6, if $\Lambda$ is a constant, then the result coincides with Theorem 5 of Embrechts et al. (2020) for $n=2$, where the semi-explicit formula is offered. Using (4), it might be difficult to find more explicit expression for the inf-convolution of $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}$ and $\rho_{g}^{\mathbb{Q}_{2}}$$\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}$.
@@ 5 Inf-convolution of $\Lambda\mathrm{VaR}^{+}$ and a monotone risk measure
The inf-convolution of multiple $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs is a very challenging problem even if we suppose the beliefs are homogeneous; see the discussion for homogeneous beliefs in Liu (2025). One important feature of $\Lambda\mathrm{VaR}^{+}$ is to capture the tail risk for decreasing $\Lambda$; see Frittelli et al. (2014) and Hitaj et al. (2018). In this section, we focus on the inf-convolution of $\Lambda\mathrm{VaR}^{+}$ and a monotone risk measure $\rho$ under $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$, which includes the inf-convolution of two $\Lambda\mathrm{VaR}^{+}$ as a special case. For any $\Lambda\in\mathcal{H}$, we denote $\Lambda_{x}(z)=\inf_{x\leqslant t\leqslant z}\Lambda(t),~z\geqslant x$. Then $\Lambda_{x}(z)$ is decreasing and right-continuous for $z\geqslant x$. If $\Lambda$ is decreasing and right-continuous, then $\Lambda_{x}(z)=\Lambda(z),~~z\geqslant x$; and if $\Lambda$ is increasing, then $\Lambda_{x}(z)=\Lambda(x),~~z\geqslant x$. The following result is valid for general $\Lambda\in\mathcal{H}$. We say $\rho$ is continuous from above on $\mathcal{X}$ if $X_{n}\downarrow X$ implies $\lim_{n\to\infty}\rho(X_{n})=\rho(X)$ for $X_{n},X\in\mathcal{X}$.
@@
Suppose $\Lambda\in\mathcal{H}$ with $0<\lambda^{-}\leqslant\lambda^{+}<1$, and $\rho$ is monotone and law-invariant under $\mathbb{Q}_{2}$. If either (i) $\mathcal{X}=L^{\infty}$ or (ii) $\mathcal{X}$ is unbounded and $\rho$ is continuous from above, then we have
@@
Moreover, for case (i), the optimal allocation of the inf-convolution exists if and only if the minimizer of (10) exists. For both cases (i) and (ii), if $(x^{*},y^{*},U^{*})$ is the minimizer, then one optimal allocation is given by
@@
In the case of $\rho=\Lambda_{1}\mathrm{VaR}^{+}$ for $\Lambda_{1}\in\mathcal{H}$ with $0<\lambda_{1}^{-}\leqslant\lambda_{1}^{+}<1$, it follows from Theorem 6 that the optimal risk allocation for risk sharing between $\Lambda\mathrm{VaR}^{+}$ and $\Lambda_{1}\mathrm{VaR}^{+}$ under $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$ has the form of (11), which is neither comonotonic nor pairwise countermonotonic. This means that the optimal risk sharing for multiple $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs is qualitatively different from the cases with convex risk measures or quantile-based risk measures in the literature. The expression of the inf-convolution in (10) is an optimization problem involving three parameters $(x,y,U)$, where $U\sim U[0,1]$ under $\mathbb{Q}_{1}$. Note that the distribution of $U$ under $\mathbb{Q}_{2}$ depends on the relation between $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$. Moreover, the expression in (10) and also the optimal allocation in (11) show that shape of $\Lambda$ function plays an important role in both the expression of the inf-convolution and the form of the optimal allocations, which is very different from the results of $\Lambda\mathrm{VaR}$ displayed in Sections 3 and 4.
For case (ii) of Theorem 6, we suppose that $\rho$ is continuous from above. In fact, many risk functionals satisfy this assumption, such as distortion risk measures $\rho_{g}$ with continuous $g$ (including $\mathrm{ES}_{\alpha}$ as a special case), expected utility $\mathbb{E}_{u}$ for continuous and increasing $u$, and the right quantiles. Later, we will define an alternative Lambda-$\mathrm{VaR}$ measure that also satisfies this assumption.
@@
We denote $X\leqslant_{\mathrm{i}cx}Y$ if $\mathbb{E}(f(X))\leqslant\mathbb{E}(f(Y))$ for all increasing and convex function $f$. A mapping $\rho:\mathcal{X}\to\mathbb{R}$ is said to be SSD-consistent if $X\leqslant_{\mathrm{i}cx}Y$ implies $\rho(X)\leqslant\rho(Y)$. Note that here $\mathcal{X}\subset L^{1}$.
Corollary 7.
Suppose $\mathcal{X}\subset L^{1}$. If $\mathbb{Q}_{1}=\mathbb{Q}_{2}=\mathbb{Q}$ and $\rho$ is SSD-consistent risk measure, then under the assumption of Theorem 6, we have
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\left\{x+\rho^{\mathbb{Q}}\left(X-\Lambda_{x,y}^{-1}(U_{X}^{\mathbb{Q}})\right)\right\}.$
Note that the above corollary coincides with Theorem 5 of Liu (2025). We next see an example that $\rho$ is not SSD-consistent. For $\Lambda:\mathbb{R}\to[0,1]$ and an increasing $f:\mathbb{R}\to[0,1]$, define
$q_{\Lambda}^{+}(f)=\inf\{x\in\mathbb{R}:f(x)>1-\Lambda(x)\},$
where $\inf\emptyset=\infty$. Note that if $\Lambda$ is a constant and $f$ is a distribution function, then $q_{\Lambda}^{+}(f)$ is the right quantile. We refer to Bellini and Peri (2022) and Han and Liu (2025) for the properties and applications of $q_{\Lambda}^{+}$. Define an alternative version of $\Lambda\mathrm{VaR}$ by $\overline{\Lambda\mathrm{VaR}}(X)=q_{\Lambda}^{+}(F_{X})$ for $X\in\mathcal{X}$. One can easily show that $\overline{\Lambda\mathrm{VaR}}$ is continuous from above if $\Lambda\in\mathcal{H}$. Then applying Theorem 6, we have the following result.
Corollary 8.
Suppose $\mathbb{Q}_{1}=\mathbb{Q}_{2}=\mathbb{Q}$, $\Lambda\in\mathcal{H}$ with $0<\lambda^{-}\leqslant\lambda^{+}<1$ is strictly decreasing and $\Lambda_{1}\in\mathcal{H}$ with $0<\lambda_{1}^{-}\leqslant\lambda_{1}^{+}<1$. Then we have, for $X\in\mathcal{X}$ with continuous $F_{X}^{\mathbb{Q},-1}$ over $(0,1)$,
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}}\mathop{\square}\displaylimits\overline{\Lambda_{1}\mathrm{VaR}}^{\mathbb{Q}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\left\{x+q_{\Lambda_{1}}^{+}(l_{x,y}^{-1})\right\},$
where $l_{x,y}(t)=\sup_{s\in(0,t]}\{F_{X}^{\mathbb{Q},-1}(s)-\Lambda_{x,y}^{-1}(1-t+s)\},~t\in(0,1)$ and $l_{x,y}^{-1}(z)=\inf\{w\in(0,1):l_{x,y}(w)\geqslant z\},~z\in\mathbb{R}$ with $\inf\emptyset=1$.
Corollary 8 requires that $F_{X}^{\mathbb{Q},-1}$ is continuous over $(0,1)$, which is a technical assumption for the computation of $\inf_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\Lambda_{1}\mathrm{VaR}}^{\mathbb{Q}}\left(X-\Lambda_{x,y}^{-1}(U)\right)$. Our computation may not be correct if this assumption is violated.
@@
For $x_{1}\in\mathbb{R}$ and $0<\lambda_{1}<\lambda_{2}<1$, let $\Lambda(z)=(1-\lambda_{1})\mathds{1}_{\{z<x_{1}\}}+(1-\lambda_{2})\mathds{1}_{\{z\geqslant x_{1}\}}$. Under the assumption of Theorem 6, we have
@@
where $\mathcal{B}_{\lambda_{1},\lambda_{2}}=\{(B_{1},B_{2}):\mathbb{Q}_{1}(B_{1})=\lambda_{1},\mathbb{Q}_{1}(B_{2})=\lambda_{2}-\lambda_{1},B_{1}\cap B_{2}=\emptyset\}$.
Note that the optimization problem in (12) is much easier than that in (10) because it only has parameters $(x,y,B_{1},B_{2})$. To illustrate our Proposition 6, we next see the example of $\rho^{\mathbb{Q}_{2}}=\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}$. In the following corollary, we suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|D_{i})$ for some $D_{i}\in\mathcal{F},~i=1,2$, and $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|D_{1}\cap D_{2})$ with $\mathbb{P}(D_{1}\cap D_{2})>0$.
Corollary 9.
For $x_{1}\in\mathbb{R}$ and $0<\lambda_{1}<\lambda_{2}<1$, let $\Lambda(z)=(1-\lambda_{1})\mathds{1}_{\{z<x_{1}\}}+(1-\lambda_{2})\mathds{1}_{\{z\geqslant x_{1}\}}$. Then we have
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)=\left\{\begin{array}[]{cc}\frac{1}{\alpha}\frac{\mathbb{P}(D_{1}\cap D_{2})}{\mathbb{P}(D_{2})}\int_{\alpha_{0}}^{\alpha_{1}}\mathrm{VaR}_{t}^{\mathbb{Q}}(X)\mathrm{d}t,&\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})\geqslant\alpha\mathbb{P}(D_{2})\\ -\infty,&\text{otherwise}\end{array}\right$.,
where $\alpha_{0}=1-\frac{\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})}{\mathbb{P}(D_{1}\cap D_{2})}$ and $\alpha_{1}=1-\frac{\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})-\alpha\mathbb{P}(D_{2})}{\mathbb{P}(D_{1}\cap D_{2})}$.
@@ 6 Conclusion
In this paper, we consider the inf-convolution of multiple $\Lambda\mathrm{VaR}$ under heterogeneous beliefs. We obtain the general expression for the inf-convolution and the corresponding optimal allocations. To study impact of the relation of the beliefs on the inf-convolution, we discuss three different cases: i) homogeneous beliefs; ii) conditional beliefs; iii) general beliefs with two agents. For all these cases, we obtain more explicit expressions of the inf-convolution. In all above cases we demonstrate that trivial outcomes arise when both belief inconsistency and risk tolerance are high. The inf-convolution of one $\Lambda\mathrm{VaR}/\Lambda\mathrm{VaR}^{+}$ and a general risk measure with belief heterogeneity are also discussed. There are still some unsolved problems, such as what the inf-convolution of multiple $\Lambda\mathrm{VaR}$ under general heterogeneous beliefs is, and what the inf-convolution of multiple $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs is, which deserves further study.
Acknowledgments
The authors thank the editor, an associate editor, and two anonymous referees for their helpful comments.
Competing Interests
The authors declare no competing interests.
Appendix A Proof of all results in Sections 2-5